# Aligning LLMs to Be Robust Against Prompt Injection

### v1

@@ Appendix C Study of the Attack Evaluation Setting
Appendix D Rule-Following Behavior
Our work aims at defending against prompt injections. A similar but orthogonal problem is called rule-following robustness, where the attacker uses a second instruction to override the first instruction (rule) [77]. Compared to prompt injections in LLM-integrated applications where there is a strictly defined prompt and data part, rule-following attacks target at LLM chatbots where there is only one user input containing multiple instructions.
We are interested to see if models robust to prompt injections are good rule-following models. We use CyberSecEval3 [78], which contains 251 samples, each with an original instruction, followed by an additional instruction asking the model to betray the original instruction. We use the termed "prompt injection test cases" in [78], which are actually performing rule-following attacks, i.e., the samples contain no benign data (prompt injection attackers should not be able to change the data part, let along erasing the benign data).
We use the CyberSecEval3 dataset with our prompt format to feed the samples to our models, and use GPT-4-turbo to judge whether the model response follows the second instruction, as [78] recommends. We directly use LLMs trained on cleaned alpaca dataset (as in all other experiments) and evaluate them with the CyberSecEval3 data. We show the results in Table 5, from which we find strong prompt injection defenses such as StruQ and SecAlign lends little (or no) robustness to rule-following attacks. This indicates that rule-following attacks are different from prompt injection, and defending against attacks without benign data tends to be harder.
Table 5: The performance of prompt injection defense against rule-following attacks.
Model Defense CyberSecEval3 ASR ($\downarrow$)
None 32
Llama-7B StruQ 25
SecAlign 27
None 39
Mistral-7B StruQ 29
SecAlign 33
None 24
Llama3-8B StruQ 20
SecAlign 36
Table 6: SecAlign breakdown results (%) from Figure 3 plus results on combining SecAlign and StruQ (“Combined”).
Model Llama-7B Mistral-7B Llama3-8B
Defense None StruQ SecAlign Combined None StruQ SecAlign Combined None StruQ SecAlign Combined
WinRate 55.46 54.55 56.06 56.85 72.21 72.17 72.88 72.46 69.47 68.77 68.87 68.67
Ignore ASR 10 0 0 0 22 0 0 0 30 0 0 0
Completion ASR 45 0 0 0 89 4 0 0 90 0 0 0
Ignore-Completion ASR 75 0.5 0 0 70 1 0 0 89 0 0 0
Max ASR (Opt.-Free) 75 0.5 0 0 89 4 0 0 90 0 0 0
AdvPrompter ASR 60 4 1 0 72 7 0 8 95 18 0 1
GCG ASR 97 58 15 5 99 56 2 6 89 33 11 22
Max ASR (Opt.-Based) 97 58 15 5 99 56 2 6 95 33 11 22
Table 7: Comparison with baselines from Table 1. Results (%) are from Llama-7B study except that BIPIA is using Vicuna-7B.
@@ Appendix D Rule-Following Behavior
Defense None Instructional Reminder Isolation Sandwich In-Context BIPIA StruQ SecAlign
Ignore ASR 10 22 20 5 3 1 1 0 0
Ignore ASR (in R) 39 47 50 39 27 22 5 0 0
Completion ASR 45 58 62 53 16 25 4 5 0
Completion ASR (in R) 71 84 75 70 34 53 4 5 0
Ignore-Completion ASR 75 78 79 73 38 45 7 0 0
Ignore-Completion ASR (in R) 84 87 83 80 44 50 7 0 0.5
Max ASR Opt.-Free 75 78 79 73 38 45 7 5 0
Max ASR Opt.-Free (in R) 84 87 83 80 44 60 7 5 0.5
Table 8: The injected word posts little influence on the evaluation results and conclusion. Results (%) are from Llama-7B study.
Hacked Hello Word Serendipity
Defense None StruQ SecAlign None StruQ SecAlign None StruQ SecAlign
Ignore ASR 10 0 0 3 0 0 28 0.5 0
Ignore ASR (in R) 39 0 0 30 1 0.5 55 3 1
Completion ASR 45 5 0 35 0 0 88 1 0
Completion ASR (in R) 71 5 0 91 1 0.5 92 1 0.5
Ignore-Completion ASR 75 0 0 73 0 0 86 1 0
Ignore-Completion ASR (in R) 84 0 0.5 85 1 0.5 91 2 0
Max ASR Opt.-Free 75 5 0 73 0 0 88 1 0
Max ASR Opt.-Free (in R) 84 5 0.5 91 1 0.5 92 3 1

### v2

@@ 4.4. Evaluation Setting Studies
Utility benchmark
We run more utility benchmarks (MMLU [Hendrycks et al., 2020], Winogrande [Sakaguchi et al., 2021], AGIEval [Zhong et al., 2023], and CommonSenseQA [Talmor et al., 2018]) on Mistral-7B and Llama3-8B to check the model’s function outside the AlpacaEval2 benchmark presented in the main experiments. Our results are presented in Table 5. In most benchmarks, SecAlign suffers from no utility score decrease. For MMLU that mostly evaluates the base model’s knowledge, the loss is 2% to 3%.
Table 5. Results on more utility benchmarks
Model Mistral-7B Llama3-8B
Defense None SecAlign None SecAlign
MMLU (%, $\uparrow$) 62.7 59.5 65.3 63.1
Winogrande (%, $\uparrow$) 77.8 77.7 77.5 77.2
AGIEval (%, $\uparrow$) 25.8 25.2 33.1 30.3
CommonSenseQA (%, $\uparrow$) 70.9 70.9 78.2 78.3
4.5. Ablation Studies
SecAlign using different preference optimization algorithms
The preference optimization algorithm is a central component in our defense. Though our contribution is not a new preference optimization technique, and the choice of it is orthogonal to SecAlign, we study the performance of SecAlign using different preference optimization besides the default DPO [Rafailov et al., 2024]. KTO [Ethayarajh et al., 2024] uses human-aware losses that maximize the generation utility instead of maximizing the log-likelihood of preferences, and is claimed to surpass DPO especially under data imbalance. ORPO [Hong et al., 2024] slightly penalizes the undesirable response in SFT to align the LLM without using additional post-SFT training, but we implement it after our SFT to align the evaluation setting with other results. We tune the leaning rates of DPO, KTO, and ORPO separately to be $[2,0.8,6.4]\times 10^{-4}$ respectively, and their $\beta$ are all 0.1. As in Table 6, all three methods exhibit similar utility performance. For security, KTO achieves the best results in our isolated experiment, albeit at the cost of a significantly increased runtime. ORPO is slightly faster but suffers from a doubled ASR. DPO emerges as the optimal balance between efficiency and performance.
Table 6. Ablation study of preference optimization algorithms in SecAlign on Llama-7B using 4 80G A100s.
[…]
DPO [Rafailov et al., 2024] 56.06 15 2 $\times$ 4
ORPO [Hong et al., 2024] 54.75 34 1.5 $\times$ 4
KTO [Ethayarajh et al., 2024] 55.84 9 10 $\times$ 4
SecAlign using different dataset sizes
SecAlign’s preference dataset effortlessly uses human-written instructions and responses from a benign SFT dataset. But the collection of SFT datasets is typically labor-intensive, especially if a diverse set of high-quality samples is needed. Consequently, a natural question to ask is whether the performance of SecAlign strongly depends on having access to a large amount of diverse SFT samples. To study this aspect, we analyze the performance when using different proportions of the training samples. We sub-sample the SFT dataset without changing the ratio of samples with a data part (those we could apply a prompt injection to). We use those datasets to perform StruQ and the first SFT step of SecAlign, then build the preference dataset using a sub-sampled SFT dataset. In this way, the number of samples seen in StruQ and SecAlign are always the same. We plot the trend in Fig. 6. Both utility and security improve as we add more training samples. SecAlign consistently maintains an ASR that is half of that observed with StruQ across different dataset portions, achieving satisfactory ASR (lower than StruQ on all samples) even with only 20% of the original samples. SecAlign demonstrates marginally higher utility when using ¿50% samples, indicating its potential when the dataset size is very large. This result shows that SecAlign can achieve a strong defense performance even under limited SFT data.
SecAlign using different learning rates
As fine-tuning LLMs involves training large neural networks, it is pertinent to examine the sensitivity of our methods to different hyperparameter choices, with the learning rate being one of the most critical. In Fig. 6, we report performance metrics across various learning rates. Intuitively, this hyperparameter noticeably impacts SecAlign. Nevertheless, various choices within a reasonable range surpass the best-performing StruQ. Additionally, SecAlign training leads to stable performance, leading to negligible error bars on utility and security as in Fig. 6 at the optimal learning rate.
Figure 6. Left: The utility (AlpacaEval2 WinRate) and security (ASR) when using different proportions of training samples. Even using 20% of the samples, SecAlign enjoys much lower ASR v.s. StruQ using all samples. Right: SecAlign enjoys equivalent utility (AlpacaEval2 WinRate) and much better security (ASR) v.s. StruQ even when tuning DPO learning rate extensively from $6\times 10^{-5}$ to $2.6\times 10^{-4}$. SecAlign is also robust to randomness in training: the two boxes in the optimal learning rate of $2\times 10^{-4}$ indicate small error bars calculated in five random runs.
5. Related Work
LLM-integrated applications.
LLMs have demonstrated remarkable success across a variety of tasks, including question-answering [Wei et al., 2022], machine translation [Zhu et al., 2023a], and summarization [Zhang et al., 2023], garnering significant attention from both academia and industry. This superiority in natural language understanding has facilitated the integration of LLMs into numerous applications, enabling the creation of task-specific models deployable via APIs [OpenAI, 2024b, Anthropic, 2023]. Recent advancements have further expanded the capabilities of LLMs, allowing for the development of AI agents capable of reasoning and planning to address complex real-world challenges, potentially leveraging third-party tools [Schick et al., 2024, Patil et al., 2023, OpenAI, 2024c]. Since AI agents interact with third-party tools containing potential unsafe data [Debenedetti et al., 2024], this wide application of LLMs introduces new risks to building a safe LLM system.
Prompt injection attacks.
Prompt injection is an emerging threat to LLM in systems [Branch et al., 2022, Perez and Ribeiro, 2022, Greshake et al., 2023, Liu et al., 2024, Toyer et al., 2024, Yu et al., 2023a, Yip et al., 2023] where an untrusted user deliberately supplies an additional instruction to manipulate the LLM functionality. Prompt injections could be categorized as direct prompt injections [Perez and Ribeiro, 2022] if the user directly types the malicious data, and indirect prompt injections [Greshake et al., 2023] if the injected data comes from an external content, e.g., a web page. Prompt injection attacks bear a conceptual similarity to traditional injection attacks in computer security. For example, in SQL injection, attackers exploit vulnerabilities by embedding malicious code into input fields, thereby manipulating SQL queries to access or alter database information [Halfond et al., 2006]. Similarly, UNIX command injection involves attackers inserting harmful commands into input fields to execute unauthorized actions on a server [Zhong et al., 2024].
Other threats to LLMs.
Alongside prompt injection, another area of LLM security research is jailbreaking attacks [Mazeika et al., 2024], which input one malicious instruction (without any data) to elicit toxic, offensive, or inappropriate outputs. Note that jailbreaking is distinct from prompt injection, where the instruction (from the system designer) is always benign and the attacker injects a prompt in the data but cannot manipulate the whole LLM input. That is, prompt injection involves a trusted system designer (providing an instruction) and an untrusted user (providing a data), but jailbreaks only involve an untrusted user (providing an instruction). Researchers have studied other attacks on LLMs, including data extraction [Carlini et al., 2021, Yu et al., 2023b, Nasr et al., 2023, Lukas et al., 2023, Li et al., 2023b] (recovering training data), membership inference attacks [Mattern et al., 2023, Duan et al., 2024] (deciding whether an existing data is in the training set), and adversarial attacks (decrease LLM’s performance) [Zhu et al., 2023b, Kandpal et al., 2023, Wang et al., 2023]. Those attacks target different LLM vulnerabilities, e.g., failure to follow prioritized instructions (prompt injections), failure to reject offensive outputs (jailbreaks), failure to provide diverse outputs than in the dataset (privacy attacks), etc. Thus, their defenses vary significantly, e.g., defenses against prompt injections separate instruction and input, while defenses against jailbreaks reject toxic inputs. However, the optimizer to realize those different attacks could be shared, as all attackers are optimizing the LLM input to elicit some specific outputs. In this work, we adapt the original jailbreaking attacks GCG [Zou et al., 2023] and AdvPrompter [Paulus et al., 2024] to do prompt injections. This could be done by simply changing the input and target output strings.
LLM alignment.
Reinforcement Learning from Human Feedback (RLHF) has emerged as a pivotal methodology for training LLMs [Ouyang et al., 2022, Kaufmann et al., 2023], allowing LLMs to align model outputs with human values and preferences, thereby ensuring more reliable, safe, and contextually appropriate responses. Within RLHF, two primary paradigms have been explored: online and offline RLHF. Offline RLHF relies on fixed, pre-collected datasets of human judgments to train a policy for LLMs. A notable example includes DPO [Rafailov et al., 2024], which we use in SecAlign. In contrast, online RLHF allows for the adaptive collection of additional preference data, either through a reward model or direct human feedback, to improve alignment. Such methods are inspired by REINFORCE [Williams, 1992] and its variants [Schulman et al., 2017]. More recently, hybrid approaches have been proposed, combining online and offline RLHF to leverage their respective strengths [Dong et al., 2024].
6. Conclusion and Discussions
We present SecAlign, a SOTA fine-tuning-based defense for securing LLMs against prompt injection using alignment. The main advantages of SecAlign are its simplicity, utility-preservation, and strong security to unseen attacks, even against optimization-based attacks. Also, through preference optimization, our work draws the connection between LLM security and alignment—two subjects that have so far been studied in separation. Our work serves as a proof-of-concept that demonstrates the efficacy of preference optimization for LLM security. Still, SecAlign has below limitations.
  • SecAlign only applies to the scenarios when the instruction part and data part are explicitly stated with clear separations (e.g., by the delimiters).
  • Our construction of desirable outputs shares a disadvantage with all existing fine-tuning-based defenses: The desirable output ignores the injected instruction in the data instead of processing it as part of the data. This may lead the LLM to ignore some imperative sentences (which may not be an injection and should be handled as data, e.g., an imperative sentence to be translated) in the data, though we do not observe this phenomenon or hurt of utility in our study. Solving it requires a careful selection of the injection (based on the benign instruction) and a specially generated desirable response that is not from the SFT dataset.
  • As a defense to AI systems, SecAlign cannot achieve 100% security yet. For stronger security in LLM-integrated applications, we suspect the need for a multi-tiered defense combining SecAlign with other techniques such as detection (e.g., Prompt Shields [202, 2024c], PromptGuard [Meta, 2024]), and input reformatting [Jain et al., 2023]. We do not regard SecAlign as a final solution to prompt injection defense.
  • In its current form, SecAlign cannot defend against attacks outside prompt injections, e.g., jailbreaks and data extraction.
Advanced fine-tuning-based defenses with SecAlign.
We apply SecAlign to a static preference dataset constructed from benign instructions and data and optimization-free injected prompts. It is plausible to further extend this idea to use optimization-based prompt injections to customize the injection to an LLM at every fine-tuning step. Applying the above idea is computationally infeasible with existing techniques. Prompt optimization remains a difficult problem due to the discrete nature of tokens. GCG, arguably the most effective optimization method right now, is too costly to run as an inner optimization loop inside SecAlign fine-tuning (estimated thousands of GPU hours are needed even for the toy Alpaca dataset). Future work on more efficient prompt optimization techniques may enable optimization-based injections in training.
Securing LLMs in real-world systems.
Securing against multi-modal prompt injections.
So far we have focused on text-only LLMs. Frontier LLMs such as GPT-4o and Gemini Pro Vision have additional input modalities such as image and/or speech, providing additional avenues for prompt injection attacks. Since these models are typically aligned using multi-modal instruction tuning, we may be able to extend SecAlign to handle protection against prompt injection in these additional input modalities [Willison, 2023b]. The new challenge here is the much easier attacks in continuous input domains (e.g., image and speech), making the attack more powerful compared to text-only prompt injection [Carlini et al., 2024]. Thus, we believe it is a new and important problem to study prompt injection defenses in these modalities.
Acknowledgments
This work is supported by Meta-BAIR Commons (Year 6). We are grateful for insightful discussions and comments from Chawin Sitawarin, Raluca Ada Popa, and anonymous reviewers.
Appendix
  • Fig. 7: results on more different-sized base models than Fig. 4.
  • Table 7: breakdown numbers in Fig. 3 and Fig. 4 (SecAlign is a SOTA fine-tuning-based defense). For optimization-free attacks, Ignore attack is the weakest, with Completion or Completion-Ignore being stronger. For optimization-based attacks, GCG is strongest tested attack on LLMs SFTed by us, and AdvPrompter turns out the be stronger on industry-tuned (Instruct) models.
  • Table 8: breakdown numbers in Table 2 (SecAlign significantly outperforms existing prompting-based defenses).
  • Table 9: breakdown numbers in Table 3 (The injection word and attack success criteria do not affect our claim).
Figure 7. Extending Fig. 4 to 6B and 13B base models.
Table 7. SecAlign is a SOTA fine-tuning-based defense: breakdown numbers from Fig. 3 and Fig. 4
Model Mistral-7B-Instruct Llama3-8B-Instruct Llama-7B Mistral-7B Llama3-8B
Defense None StruQ SecAlign None StruQ SecAlign None StruQ SecAlign None StruQ SecAlign None StruQ SecAlign
WinRate (%, $\uparrow$) 67.01 70.73 69.22 85.39 80.79 85.88 55.46 54.55 56.06 72.21 72.17 72.88 69.47 68.77 68.87
Ignore ASR (%, $\downarrow$) 18 0.5 0 24 0 0 10 0 0 22 0 0 30 0 0
Completion ASR (%, $\downarrow$) 59 1 0 47 0 0 45 0 0 89 4 0 90 0 0
Ignore-Completion ASR (%, $\downarrow$) 59 2 0 51 0 0 75 0.5 0 70 1 0 89 0 0
Max ASR Opt.-Free (%, $\downarrow$) 59 2 0 51 0 0 75 0.5 0 89 4 0 90 0 0
AdvPrompter ASR (%, $\downarrow$) 89 34 1 98 62 1 60 4 1 72 7 0 95 18 0
GCG ASR (%, $\downarrow$) 89 15 1 84 4 0 97 60 14 95 41 1 98 43 9
NeuralExec ASR (%, $\downarrow$) 20 16 0 63 0.5 0 2 0 0 32 2 0 34 0 0
Max ASR Opt.-Based (%, $\downarrow$) 89 34 1 98 62 1 97 60 14 95 41 1 98 43 9
Table 8. SecAlign significantly outperforms existing prompting-based defenses: breakdown numbers from Table 2.
@@ Acknowledgments
Defense Model None Instructional Reminder Isolation Sandwich In-Context SecAlign
Ignore ASR (%, $\downarrow$) Llama3-8B-Instruct 24 16 18 27 16 0.5 0
Completion ASR (%, $\downarrow$) 47 31 21 35 16 0.5 0
Ignore-Completion ASR (%, $\downarrow$) 51 38 35 50 53 0 0
Max ASR Opt.-Free (%, $\downarrow$) 51 38 35 50 55 0.5 0
Ignore ASR (%, $\downarrow$) Llama-7B 10 22 20 5 3 1 0
Completion ASR (%, $\downarrow$) 45 58 62 53 16 25 0
Ignore-Completion ASR (%, $\downarrow$) 75 78 79 73 38 45 0
Max ASR Opt.-Free (%, $\downarrow$) 75 78 79 73 38 45 0
Table 9. The injection word and attack success criteria do not affect our claim: breakdown numbers from Table 3.
Injected Word Hacked Hello Word Serendipity
Max ASR Opt.-Free w. Diff. Attack Success Criteria None StruQ SecAlign None StruQ SecAlign None StruQ SecAlign
Ignore Begin-With (default) ASR (%, $\downarrow$) 10 0 0 3 0 0 28 0.5 0
Ignore In-Response ASR (%, $\downarrow$) 39 0 0 30 1 0.5 55 3 1
Completion Begin-With (default) ASR (%, $\downarrow$) 45 5 0 35 0 0 88 1 0
Completion In-Response ASR (%, $\downarrow$) 71 5 0 91 1 0.5 92 1 0.5
Ignore-Completion Begin-With (default) ASR (%, $\downarrow$) 75 0 0 73 0 0 86 1 0
Ignore-Completion In-Response ASR (%, $\downarrow$) 84 0 0.5 85 1 0.5 91 2 0
Max Begin-With (default) ASR Opt.-Free (%, $\downarrow$) 75 5 0 73 0 0 88 1 0
Max In-Response ASR Opt.-Free (%, $\downarrow$) 84 5 0.5 91 1 0.5 92 3 1